A satellite moves in an elliptical orbit with the earth at one focus . At the perigee it's speed is v and it's distance from the centre of the earth is r . If the eccentricity of orbit is 0.5, calculate it's speed at the apogee
1
Expert's answer
2019-03-28T08:36:33-0400

satellite velocity on elliptical orbit is

"v^2=GM(\frac{2}{r}-\frac{1}{a})"

but for apogee this this equation is

"r=a(1+e)"


"v_A^2=\frac{GM}{a} \times\frac{(1-e)}{(1+e)}"

for perigee

"r=a(1-e)"

"v_P^2=\frac{GM}{a} \times\frac{(1+e)}{(1-e)}"

Thus

"\frac{v_A^2}{v_P^2}=\frac{\frac{GM}{a} \times\frac{(1-e)}{(1+e)}}{\frac{GM}{a} \times\frac{(1+e)}{(1-e)}}"

we have

"v_A=\frac{(1-e)}{(1+e)}\times v_P"

finally using initial conditions, we have

"v_A=\frac{1}{3}\times v"

satellite velocity at apogee is three times less than perigee